OpenStudy (anonymous):

Which of the following numbers would you multiply -3/4 by to equal 1?

OpenStudy (anonymous):

multiply by negative 4/3

OpenStudy (anonymous):

two negative equal a positive. Multiplying it by its inverse(a/b=b/a) 4/3 cancels each other out to get 1

OpenStudy (anonymous):

-3/4 * -4/3= 12/12 12/12= 1 negative * negative= positive hope I helped